40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 16 00:43:36 2011, Koji, Update, SUS, Vertex SUS Binary Output Boxes were turned off / need investigation 
    Reply  Thu Jun 16 08:45:14 2011, steve, Update, SUS, Vertex SUS Binary Output Boxes removed P1070894.JPG
       Reply  Thu Jun 16 23:19:09 2011, Koji, Update, SUS, Vertex SUS Binary Output Boxes removed 
Message ID: 4829     Entry time: Thu Jun 16 23:19:09 2011     In reply to: 4828
Author: Koji 
Type: Update 
Category: SUS 
Subject: Vertex SUS Binary Output Boxes removed 

[Jamie, Koji]

- We found the reason why some of the LEDs had no light. It was because the LEDs were blown as they were directly connected to the power supply.
The LEDs are presumably designed to be connected to a 5V supply (with internal current-limiting resistor of ~500Ohm). The too much current
with the 15V (~30mA) made the LED blown, or the life-time of them shorter.

- Jamie removed all of the BO modules and I put 800Ohm additional resister such that the resultant current is to be 12mA.
The LEDs were tested and are fine now.

- The four BO boxes for C1SUS were restored on the rack. I personally got confused what should be connected where
even though I had labeled for BO0 and BO1. I just have connected CH1-16 for BO0. The power supplies have been connected only to BO0 and BO1.

- I tested the whitening of PRM UL sensor by exciting PRM UL sensor. The transfer function told us that the pendulum response can be seen
up to 10-15Hz. When the whitening is on, I could see the change of the transfer function in that freq band. This is good.
So the main reason why I could not see theis was that the power supply for the BOs were not turned on.

- I suppose Jamie/Joe will restore all of the BO boxes on the racks tomorrow. I am going to make a test script for checking the PD whitenings.

ELOG V3.1.3-